# GIVT: Generative Infinite-Vocabulary Transformers

### v2

@@ Abstract
We introduce generative infinite-vocabulary transformers (GIVT) which generate vector sequences with real-valued entries, instead of discrete tokens from a finite vocabulary. To this end, we propose two surprisingly simple modifications to decoder-only transformers: 1) at the input, we replace the finite-vocabulary lookup table with a linear projection of the input vectors; and 2) at the output, we replace the logits prediction (usually mapped to a categorical distribution) with the parameters of a multivariate Gaussian mixture model. Inspired by the image-generation paradigm of VQ-GAN and MaskGIT, where transformers are used to model the discrete latent sequences of a VQ-VAE, we use GIVT to model the unquantized real-valued latent sequences of a VAE. When applying GIVT to class-conditional image generation with iterative masked modeling, we show competitive results with MaskGIT, while our approach outperforms both VQ-GAN and MaskGIT when using it for causal modeling. Finally, we obtain competitive results outside of image generation when applying our approach to panoptic segmentation and depth estimation with a VAE-based variant of the UViM $*{}^{*}$ork done as a Student Researcher at Google DeepMind.
${}^{\circ}$ignificant technical contributions.
@@ 1 Introduction
After becoming the dominant architecture in natural language processing shortly after their introduction, Transformers [67] have also recently become very popular in computer vision [16, 57, 37]. Dosovitskiy et al. [16] showed that by splitting images into sequences of patches, linearly embedding those patches, and then feeding the resulting sequence of features to a transformer encoder leads to powerful image classifiers that outperform CNN-based architectures at large model and data scale. This strategy is now standard for many discriminative vision task including classification [16], detection [37], and segmentation [57].
Figure 1: We compare the standard discrete-token generative transformer (left) to our continuous, infinite-vocabulary variant (GIVT, right), using the same decoder-only architecture. At the input, GIVT linearly embeds a sequence of real-valued vectors instead of discrete tokens via lookup. At the output, GIVT predicts the parameters of a continuous distribution over real-valued vectors rather than a categorical distribution over a finite vocabulary.
@@
Figure 2: Selected samples from the GIVT-Causal for 10 ImageNet classes (130, 130, 138, 144, 933, 145, 360, 207, 829, 248).
However, mapping an image to a sequence of (unquantized) feature vectors is not suitable for transformer-based image generation since generative transformers decoders consume and predict discrete tokens from some fixed, finite vocabulary. Such a structure naturally fits natural language, for which decoder-only models enable powerful sequential generative modeling and efficient training via teacher forcing [67, 47]. To harness these capabilities for images, recent works [18, 6, 7, 36, 43] have employed a two-stage approach which first learns a Vector-Quantized Variational Autoencoder (VQ-VAE) [66] to map images to a sequence of discrete tokens, and then learns a transformer decoder to model the latent discrete-token distribution. Another advantage of such VQ-VAE-based image tokenization is that it enables interleaved multimodal generative models, simply by concatenating the vocabularies of the different modalities including text and images [1, 26, 2].
@@
While this two-stage approach was successfully applied to image and multimodal generation, it has several issues. The vocabulary size in VQ-VAE determines how much information can be stored in the latent code sequence, and how much of the visual modeling is done by the VQ-VAE decoder. Having a small vocabulary can make the latent modeling easier but also makes the latent code less informative, which prevents control of the low-level details in image generation, and impacts quality when using the tokens for dense prediction [30, 39] or low-level discriminative tasks [1, 26]. To mitigate this issue, the vocabulary size can be increased, which, however, can lead to low vocabulary utilization [43] so that high-fidelity VQ-VAE setups rely on a range of advanced techniques, such as entropy losses [6] or codebook-splitting [30]. Furthermore, large vocabularies lead to correspondingly large embedding matrices consuming a lot of memory, which can be an issue particularly in multimodal contexts with vocabularies combined across modalities.
@@
To avoid these issues, we propose a modification to decoder-only transformers which removes the need for discrete tokens and therefore fixed, finite vocabularies. Specifically, we propose—to our knowledge for the first time—a generative transformer decoder which operates on sequences of real-valued vectors. As real-valued vectors can be seen as an infinite vocabulary, we call this a Generative Infinite-Vocabulary Transformer (GIVT). We make two minimal changes compared to the standard transformer decoder architecture [67, 47], see Fig. 1: 1) at the input, rather than using a sequence of discrete tokens to look up a finite vocabulary of embeddings, we linearly embed a sequence of real-valued vectors; and 2) at the output, rather than predicting the parameters of a categorical distribution over a finite vocabulary (via logits), we predict the parameters of a continuous distribution over real-valued vectors. We train this model in the same way as standard transformer decoders: with a causal attention mask and teacher forcing [67], and alternatively also explore fast progressive masked-bidirectional-modelling as in MaskGIT [12, 6, 7].
While GIVT can, in principle, be applied to any sequence of feature vectors, it can be challenging to directly model sequences that are prohibitively long or follow a complex distribution (e.g., the sequence of RGB pixels obtained by flattening a high-resolution image). Therefore, similar to the two-stage approach with VQ-VAEs and analogous the two-stage approach of latent-diffusion models [50, 46], we first learn a lower-dimensional latent space with a Gaussian-prior VAE, and then model it with GIVT.
Additionally, we port various inference schemes from the sequence-modeling literature (e.g., temperature sampling, classifier-free guidance). Surprisingly, this leads to a model which is comparable to or better than VQ-based approaches, by purely relying on real-valued tokens.
@@
    We show that GIVT matches or outperforms the standard discrete-token transformer decoder on both image generation and dense prediction tasks like semantic segmentation and monocular depth estimation (via UViM [30]).
@@
    We derive variants of standard sampling techniques for the continuous case, such as temperature sampling, beam search, and classifier-free guidance (CFG, [23]), and demonstrate their effectiveness.
@@
    We study the relationship between the degree of VAE latent-space regularization (via KL-term weighting [27, 22]) and the resulting properties of GIVT.
We emphasize that the VAE and GIVT training only relies on standard techniques from the deep-learning toolbox, and not the advanced training techniques of the VQ-VAE literature like auxiliary losses [66, 6] on the latent representation, codebook reinitialization [34], or dedicated optimization algorithms [30, 25].
@@ 2 Related work
VQ-VAE for visual data tokenization  Following the success of pixel-space autoregressive modeling [65, 53, 40, 8] for image generation, moving the autorgressive modeling to the latent space of VQ-VAEs [66, 49] emerged as a more efficient alternative. The use of GANs and perceptual losses for VQ-VAE training as well as modern causal [18, 72, 68] and masked [6, 7, 36] transformers for latent modeling led to substantial quality improvements. Another active area leveraging VQ-VAEs is interleaved multimodal generative modeling of images and text [1, 26, 2]. Further, VQ-VAEs are a popular choice to tokenize the label space of dense prediction vision tasks [30, 39]. Finally, some language-inspired techniques for self-supervised learning from images rely on VQ-VAE representations [3, 70, 36].
@@
VAEs with learned priors  A rich body of literature studies improving VAEs with learned priors: Inverse autoregressive flows emerged as a popular choice [28, 9]. Other approaches are based on normalizing flows [63] or a mixture of variational posteriors with pseudo-inputs [60]. For VAEs with discrete (non-VQ) latents, learned priors based on Restricted Boltzmann Machines were studied [64, 51].
@@ 3.1 VAE training
We first train a continuous $\beta$-VAE [22] with Gaussian encoder and prior as originally proposed by [27]. Given an input image $x$, the encoder $E$ predicts mean $\mu$, and variance $\sigma$ of a multivariate normal distribution with diagonal covariance matrix, and samples a representation $z$ from $\mathcal{N}(\mu,\sigma)$ using the reparameterization trick [27]. The VAE decoder then maps the latent sequence back to an image. Since we use a Gaussian encoder distribution, the KL-term in the evidence lower bound (ELBO) [27] can be computed in closed form as described in [27, Sec. F.1]. As for the reconstruction/likelihood term in the ELBO, we rely on a mixture of MSE, perceptual loss and GAN loss for image generation following [6, 18], or the categorical cross-entropy for dense prediction tasks [30]. Our encoder spatially-downsamples $x$, whereby we obtain a $z$ with spatial dimensions $h\times w$ and feature dimension $d$, with $h{=}\lceil H/16\rceil,w{=}\lceil W/16\rceil$, given a $H{\times}W$ input $x$.
@@
DiT-XL/2-G [46] CFG $=1.50$ 250 $\numprint{2.27}$ $\numprint{0.83}$ $\numprint{0.57}$
Sequence VQGAN [18] Top-k $=$ Mixed ${}^{\star}$ $\numprint{17.04}$
Models VQGAN [18] Top-k $=600$, CG $=0.05$ ${}^{\star}$ $\numprint{5.20}$

### v4

@@ Abstract
We introduce Generative Infinite-Vocabulary Transformers
(GIVT) which generate vector sequences with real-valued entries, instead of discrete tokens from a finite vocabulary. To this end, we propose two surprisingly simple modifications to decoder-only transformers: 1) at the input, we replace the finite-vocabulary lookup table with a linear projection of the input vectors; and 2) at the output, we replace the logits prediction (usually mapped to a categorical distribution) with the parameters of a multivariate Gaussian mixture model. Inspired by the image-generation paradigm of VQ-GAN and MaskGIT, where transformers are used to model the discrete latent sequences of a VQ-VAE, we use GIVT to model the unquantized real-valued latent sequences of a $\beta$-VAE. In class-conditional image generation GIVT outperforms VQ-GAN (and improved variants thereof) as well as MaskGIT, and achieves performance competitive with recent latent diffusion models. Finally, we obtain strong results outside of image generation when applying GIVT to panoptic segmentation and depth estimation with a VAE variant of the UViM framework. Work done as Student Researcher at GDM. Significant technical contributions. Code and model checkpoints: https://github.com/google-research/big_vision.
Keywords:
Image generation Latent sequence modeling Soft tokens
@@ 1 Introduction
After becoming the dominant architecture in natural language processing shortly after their introduction, Transformers [72] have also recently become very popular in computer vision [18, 63, 40]. Dosovitskiy et al. [18] showed that by splitting images into sequences of patches, linearly embedding those patches, and then feeding the resulting sequence of features to a transformer encoder leads to powerful image classifiers that outperform CNN-based architectures at large model and data scale. This strategy is now standard for many discriminative vision task including classification [18], detection [40], and segmentation [63]. It is less obvious how to apply generative transformer decoders to image generation since they were designed to consume and predict discrete tokens from some fixed, finite vocabulary. Such a structure naturally fits natural language, for which decoder-only models enable powerful sequential generative modeling and efficient training [72, 52].
@@
Figure 1: Selected $512\times 512$ samples from GIVT-Causal-L for 10 ImageNet classes (130, 130, 138, 144, 933, 145, 360, 207, 829, 248).
@@
To harness these capabilities for images, recent works [54, 20, 7, 6, 39, 46] have employed a two-stage approach which first trains a Vector-Quantized Variational Autoencoder (VQ-VAE) [49] to map images to a sequence of discrete tokens, and then trains a transformer decoder to model the latent discrete-token distribution. An advantage of such a VQ-VAE-based image tokenization is that it enables interleaved multimodal generative models, simply by concatenating the vocabularies of the different modalities including text and images [1, 29, 2]. However, this approach also has several issues. First, the non-continuous nature of VQ requires differentiable approximations to enable stochastic gradient-based optimization [49]. Second, a VQ-VAE with a small vocabulary can make the latent modeling easy but also makes the latent code less informative, which prevents control of the low-level details in image generation, and impacts quality when using the tokens for dense prediction [33, 42] or low-level discriminative tasks [1, 29]. A large vocabulary, on the other hand, can lead to low vocabulary utilization [46] so that high-fidelity VQ-VAE setups typically rely on a range of advanced techniques, such as entropy losses [7] or codebook-splitting [33]. Furthermore, large vocabularies lead to correspondingly large embedding matrices and hence memory consumption, which can be an issue particularly in multimodal contexts.
@@
In this work, we show—to our knowledge for the first time—how to completely remove quantization from generative transformers for visual data. Indeed, practitioners seem to agree that this would be hardly possible, since transformer decoders are strongly linked to discrete representations in many heads. Surprisingly, we not only show that simple modifications enable transformer decoders to directly generate sequences of unquantized vectors, but also that this approach leads to better image generation quality and representation learning capabilities than VQ-based approaches. We call such transformers Generative Infinite-Vocabulary Transformer (GIVT). [^1] We discuss the relation between continuous latents and infinite vocabulary in Sec. 0.A. Concretely, we make two changes compared to the standard transformer decoder architecture [72, 52], see Fig. 3: 1) at the input, rather than using a sequence of discrete tokens to look up a finite vocabulary of embeddings, GIVT linearly embeds a sequence of real-valued vectors; and 2) at the output, rather than predicting a categorical distribution over a finite vocabulary, GIVT predicts the parameters of a $d$-variate Gaussian Mixture Model (GMM). We train GIVT in the same way as standard transformer decoders: with a causal attention mask and teacher forcing [72], and alternatively also explore fast progressive masked-bidirectional-modelling as in MaskGIT [13, 7, 6].
Figure 2: We compare the standard discrete-token generative transformer (left) to our continuous, infinite-vocabulary variant (GIVT, right), using the same decoder-only architecture. At the input, GIVT linearly embeds a sequence of real-valued vectors instead of discrete tokens via lookup. At the output, GIVT predicts the parameters of a multivariate, continuous distribution rather than a categorical distribution.
Similar to the two-stage approach with VQ-VAEs and analogous the two-stage approach of latent-diffusion models [55, 51], we first learn a lower-dimensional latent space with a Gaussian-prior $\beta$-VAE [30, 24], and then model it with GIVT. We emphasize that training both $\beta$-VAE and GIVT only relies on standard techniques from the deep-learning toolbox, and not the advanced training techniques of the VQ-VAE literature like auxiliary losses [49, 7] on the latent representation, codebook reinitialization [37], or dedicated optimization algorithms [33, 27].
@@
    We show that GIVT outperforms VQGAN [55] (and follow-up variants) and MaskGIT [7] in class-conditional image generation, often by a large margin and/or at significantly lower computational cost. GIVT is also competitive with strong latent diffusion baselines, particularly at high resolution.
@@
    We derive variants of standard sampling techniques for the continuous case, such as temperature sampling, beam search, and classifier-free guidance (CFG)[25], and showcase their effectiveness.
@@
    We demonstrate that GIVT matches or outperforms prior sequential image generation models in representation learning at significantly lower computational cost.
  4. 4.
    GIVT achieves comparable performance with the VQ-based UViM approach [33] in dense prediction tasks like semantic segmentation and monocular depth estimation.
We emphasize that advances in transformer decoder-based models for visual data generation as GIVT directly benefit form advances in scaling and inference efficiency for large language models. Conversely, and unlike for diffusion models, improvements in models as ours are straight-forward to transfer to multimodal interleaved modeling [1, 29, 2] which is becoming increasingly popular.
@@ 2 Related work
VQ-VAE for visual data tokenization  Following the success of pixel-space autoregressive modeling [71, 59, 50, 43, 8] for image generation, moving the autorgressive modeling to the latent space of VQ-VAEs [49, 54] emerged as a more efficient alternative. The use of GANs and perceptual losses for VQ-VAE training as well as modern causal [20, 77, 73] and masked [7, 6, 39] transformers for latent modeling led to substantial quality improvements. Another active area leveraging VQ-VAEs is interleaved multimodal generative modeling of images and text [1, 29, 2]. Further, VQ-VAEs are a popular choice to tokenize the label space of dense prediction vision tasks [33, 42]. Finally, some language-inspired techniques for self-supervised learning from images rely on VQ-VAE representations [3, 75, 39].
@@
VAEs with learned priors  A rich body of literature studies improving VAEs with learned priors: Inverse autoregressive flows emerged as a popular choice [31, 9]. Other approaches use normalizing flows [70] or a mixture of variational posteriors with pseudo-inputs [66]. For VAEs with discrete (non-VQ) latents, learned priors based on Restricted Boltzmann Machines were studied [69, 57].
@@ 3.1 VAE training
We first train a continuous-latent $\beta$-VAE [24] with Gaussian encoder and prior as originally proposed by [30]. Given an input image $x$, the encoder $E$ predicts mean $\mu$, and covariance $\sigma$ of a multivariate normal distribution with diagonal covariance matrix, and samples a representation $z$ from $\mathcal{N}(\mu,\sigma)$ using the reparametrization trick [30]. The VAE decoder then maps the latent sequence back to an image. Since we use a Gaussian encoder distribution, the KL-term in the evidence lower bound (ELBO) [30] can be computed in closed form as described in [30, Sec. F.1]. As for the reconstruction/likelihood term in the ELBO, we rely on a mixture of MSE, perceptual loss and GAN loss for image generation following [20, 7], or the categorical cross-entropy for dense prediction tasks [33]. Our encoder spatially-downsamples $x$, whereby we obtain $z$ with spatial dimensions $h\times w$ and feature dimension $d$, with $h{=}\lceil H/16\rceil,w{=}\lceil W/16\rceil$, given a $H{\times}W$ input $x$. To compute the KL-term, the associated $\mu$ and $\sigma$ with shapes $w\times h\times d$ are flattened into $whd$ vectors.
@@
Table 1: Results on class-conditional $256{\times}256$ ImageNet, where GIVT-Causal models outperform their quantization-based counterparts at much smaller model size (VQGAN) or substantially shorter sequence length (ViT-VQGAN). We report FID as well as precision and recall (where available). We use the standard ADM evaluation suite, where FID is calculated w.r.t. the training set. +A: GIVT variants with adapter, CG: Classifier guidance acceptance rate or scale, CFG $=w$: Classifier-free guidance with weight $w$ [25], DB-CFG $=w$: Our distribution based CFG variant (Sec. 3.4), Top-k: Top-k sampling [21] (“mixed” refers to multiple $k$), $t$: Temperature sampling by scaling the predicted $\sigma$ of our models with $t$, $t_{C}$: Choice temperature for MaskGIT. Steps number of inference steps. Additional comments: Numbers obtained by us from public code, Inference uses activation caching.
@@
LDM-4 [55] 250 $\numprint{10.56}$ $\numprint{0.71}$ $\numprint{0.62}$
LDM-4-G [55] CFG $=1.5$ 250 $\numprint{3.60}$ $\numprint{0.87}$ $\numprint{0.48}$
@@
DiT-XL/2-G [51] CFG $=1.5$ 250 $\numprint{2.27}$ $\numprint{0.83}$ $\numprint{0.57}$
[…]
Modeling GIVT-MaskGIT (Ours) $t_{C}=35$ 16 $\numprint{4.64}$ $\numprint{0.85}$ $\numprint{0.49}$
GIVT-MaskGIT (Ours) $t_{C}=60$, DB-CFG $=0.1$ 16 $\numprint{4.53}$ $\numprint{0.87}$ $\numprint{0.47}$
Sequence VQGAN [20] Top-k $=$ Mixed 256 $\numprint{17.04}$
Models VQGAN [20] Top-k $=600$, CG $=0.05$ 256 $\numprint{5.20}$
ViT-VQGAN-L [76] 1024 $\numprint{4.17}$
ViT-VQGAN-L [76] CG $=0.5$ 1024 $\numprint{3.04}$
[…]
GIVT-Causal (Ours) $t=0.95$, DB-CFG $=0.4$ 256 $\numprint{3.35}$ $\numprint{0.84}$ $\numprint{0.53}$
[…]
3.2 GIVT training
We next train a GIVT to predict $p(z)$ or $p(z|c)$ (when a conditioning signal $c$ is available, e.g., in class-conditional generation). The representation $z$ is reshaped into a $hw$-length sequence of $d$-dimensional real-valued vectors (or “soft tokens”). Note how this differs from the standard VQ-VAE-based setup, where the latent transformer decoder models a $hw$-length sequence of integers denoting codebook indices. To accommodate this difference, we make two small changes to the standard transformer decoder-only architecture (see Fig. 3): We replace the embedding lookup tables at the input with a single linear layer to project from $d$ to the transformer hidden dimension. At the output, we do not predict a categorical distribution, and instead let the transformer predict the parameters of a continuous distribution. Assuming channel-wise independence of the mixture components, we model this continuous distribution with a $k$-mixture GMM. The GIVT model hence predicts $2kd+k$ parameters per soft token ($kd$ mean and $kd$ variance parameters for the mixture components, and $k$ mixture probabilities). Experimentally, we found it beneficial to normalize the mixture probabilities with a softmax activation, and the variance parameters with softplus.
We use the standard cross-entropy loss (which is equivalent to the negative log-likelihood) on the distribution $\tilde{p}$ predicted by GIVT, and minimize $\mathcal{L}_{\text{T}}=\sum_{c}\mathbb{E}_{z}\left[-\log\tilde{p}(z|c)\right]$, assuming the the classes or conditioning signal $c$ uniformly distributed (see App. 0.C.1 for details on the loss). We train two types of GIVT models, as described next.
GIVT-Causal  Here, GIVT is trained to predict every $d$-dimensional vector in the $hw$ sequence of latents conditioned on all previous vectors. Thereby, the self-attention layers are masked to be temporally causal [20, 72] (which enables sequential generation at inference time and is unrelated to causal inference). This training strategy is also called teacher forcing and is analogous to the latent modeling in VQ-GAN [20]. For class-conditional image generation we prepend a [CLS] vector to the input sequence, i.e., a learned vector for each class $c$.
GIVT-MaskGIT  As in MaskGIT [7], we mask a subset of the input sequence randomly during training and then gradually uncover the masked tokens during inference. The only changes compared to [7] are related to our real-valued tokens: since we have infinitely many tokens, there is no obvious choice to define a special mask token (when using VQ, one can just extend the vocabulary to contain special tokens, such as [MASK]). Instead, given $z$ and a mask $M$ indicating for every location whether it is masked, we first replace the locations in $z$ corresponding to $M$ with zeros (to remove information), and then embed it with a single dense layer, as above. Additionally, we concatenate one of two learned special vectors in the feature dimension, a [MASK] vector for masked locations, and a [UNMASK] vector otherwise (we half the dimension of the embedded inputs and special tokens s.t. the final hidden dimension remains unchanged).
3.3 Towards end-to-end training: Adapters
An interesting consequence of using an unquantized VAE and modeling the resulting latent sequence with a continuous rather than a categorical distribution is that the VAE and GIVT can be jointly trained or fine-tuned end-to-end (using the reparametrization trick [30]). However, this setup comes with its own set of challenges (e.g., it encompasses multiple losses which have to be balanced appropriately) and we leave it for future work. Instead, we explore a simple alternative to better match the latent distributions of the VAE and the one predicted by GIVT: We use a small invertible flow model [15, 16], or “adapter”, to map the VAE latent sequences to a new latent space of identical dimensions. We rely on a “volume preserving” additive coupling layer-based model which has a diagonal Jacobian [15]. GIVT is then trained jointly with the adapter to predict the sequences in this transformed latent space induced by the adapter (using the same loss). At inference time, samples drawn from GIVT are first processed by the inverted adapter, and then decoded to an image with the VAE decoder. Note that the adapter does not require additional losses thanks to invertibility and adds a negligible compute and model parameter overhead (less than 0.1%) compared to the GIVT model (see Sec. 4 and App. 0.B for details).
3.4 Inference
Given a VAE and GIVT trained as above, during inference we sample form GIVT either sequentially (see Fig. 3) or as in MaskGIT [7] and decode the sampled sequence into an image. We now investigate the various inference schemes for discrete transformers, and derive their continuous counterparts.
Temperature Sampling, Nucleus Sampling, Beam Search  In sequence models for text (see [26] for an overview and discussion) and VQ-GAN-based approaches, it is common to adapt and tune the sampling algorithm. We start with temperature sampling, which for discrete models adapts the softmax temperature of the categorical distributions predicted at each decoding step. For GIVT, we instead scale the covariance matrices of the predicted Gaussian distributions and call this strategy “variance scaling”. As we will see in Sec. 4, this simple change can have a significant impact on sample quality.
Nucleus sampling [26] proposes to collect the largest logits such that its cumulative probability after normalization exceeds a threshold (for example 0.8), and to sample from this reduced-support distribution. In GIVT, when predicting a single mixture, this can be approximated by truncating the predicted distributions per dimension (thereby choosing a higher-density support). This has a similar effect to variance scaling and therefore do not pursue this strategy.
Figure 4: $\beta$-VAE ablation: Interplay of KL weight $\beta$, number of channels $d$, and number of mixtures $k$ when training the VAE. Round markers show the sampling FIDs obtained with a Base-size GIVT-Causal. As $\beta$ and $k$ increase, the sampling FID improves, but the reconstruction FID also increases, limiting the best possible sampling FID.
Figure 5: Effect of different sampling strategies and model variants (GIVT-Causal-L) on sample quality. Increasing the number of mixtures $k$ and adding an adapter (+A) lead to compounding improvements. DB-CFG is the most effective sampling strategy for all model configurations.
Distribution-Based Classifier-Free Guidance  In the diffusion literature, classifier-free guidance (CFG) [25] has been employed with great success. Concretely, conditional diffusion models are trained with an additional null class $\emptyset$ to learn the unconditional data distribution. Then, during inference, the conditional log density is “moved away” from the unconditional one: given a guidance weight $w$, the updated (diffusion) score estimate is is obtained as
where $\epsilon$ estimates the gradient of the log density of the data distribution, $\epsilon(z,c)\propto\nabla_{z}\log\tilde{p}(z|c)$ (see [25, Sec. 2]). From this, we now derive a CFG variant for our GIVT, since we directly predict a density. We term this approach “Density-Based CFG” (DB-CFG). Eq. 1 can be written as
$\tilde{\epsilon}(z,c)$ $\propto(1+w)\nabla_{z}\log\tilde{p}(z|c)-w\nabla_{z}\log\tilde{p}(z|\emptyset)$
$\propto\nabla_{z}\log\left(\tilde{p}(z|c)^{1+w}\tilde{p}(z|\emptyset)^{-w}\right),$
i.e., $\tilde{\epsilon}$ estimates the log of the density $p_{\text{CFG}}(z|c)\propto\tilde{p}(z|c)^{1+w}\tilde{p}(z|\emptyset)^{-w}$ (see Fig. 6). Thus, we want to adapt our models to sample from $p_{\text{CFG}}$. We follow [25] and train GIVT with an additional null class $\emptyset$. During inference, we evaluate GIVT twice at every step, once conditional on the actual label $c$ and once conditional on $\emptyset$. To implement classifier-free guidance, we then have to sample from an unnormalized version of $p_{\text{CFG}}(z)$ derived from the two GIVT predictions. To this end, we turn to rejection sampling, which requires: 1) an unnormalized density; 2) a good proposal distribution $p^{\prime}$, that is close to the true target distribution; and 3) a scaling factor $K$ to bound the likelihood ratio between $p^{\prime}$ and the unnormalized target density.
Figure 6: Visualization of our Density-Based Classifier-Free Guidance (DB-CFG). We show the conditional and unconditional PDFs predicted by GIVT, and the resulting CFG PDF for two values of $w$. Note how the CFG distributions become more peaked. We use rejection sampling to sample from $p_{\text{CFG}}$.
The distributions we mix are GMMs and finding a good proposal distribution can be challenging. Instead, we first sample the mixture index from $\tilde{p}(z|c)$ and apply DB-CFG to the corresponding mixture components from $\tilde{p}(z|c)$ and $\tilde{p}(z)$ (the components are multivariate Gaussians with diagonal covariance). We find empirically that the unconditional components (i.e., distributions predicted using the $\emptyset$ label) tend to have larger variance than the conditional ones (as visualized in Fig. 6). It is thus sensible to pick sample proposals from $\mathcal{N}(\mu_{c},2\sigma_{c})$, where $\mu_{c},\sigma_{c}$ are the parameters predicted by GIVT when given the label $c$. We empirically find that drawing 1000 samples is enough to find at least one valid sample 99.9% of the time. For the remaining ${<}0.1$%, fall back to sampling from $\mathcal{N}(\mu_{c},\sigma_{c})$.
We emphasize that the overhead of DB-CFG is small: it requires two forward passes (per inference step) instead of one to predict the conditional and unconditional distribution. We then draw 1000 samples from those in parallel on an accelerator, which is very fast. We refer to App. 3.4 for Python code.
Figure 7: Left: Impact of DB-CFG (Sec. 3.4) and variance scaling (Sec. 3.4) on sampling FID of our class-conditional $256\times 256$ GIVT-Causal models. DB-CFG values in $[0.3,0.8]$ and variance scaling parameter $t$ in $[0.9,1.0]$ lead to low FID. Right: Average standard deviation of the GMM predicted by GIVT-Causal for class 130, averaged over 128 samples: conditional predictions have lower standard deviation; spikes can be observed when the line changes in the raster scan over the latent feature vectors.
4 Experiments
4.1 Image generation
We use ImageNet1k [56] and explore class-conditional generation (where we condition our GIVT on class labels) for 256px and 512px, and unconditional generation for 256px.
$\beta$-VAE  We closely follow the setup of MaskGIT [7]. We use their VAE architecture, built of ResBlocks (as detailed in App. 0.C; encoder and decoder have a combined 53.5M parameters), remove the VQ layer and related losses, and replace it with a linear layer predicting $\mu,\sigma$ (Sec. 3.1). We use the same weights for reconstruction, perceptual, and GAN-loss, as well as identical optimizer parameters, as in [7, 46]; we only vary the latent dimension $d$ and weight $\beta$ of the KL-term. By default, we set the token dimension to $d=16$ (i.e., the VAE predicts 16 means and variances per token) and $\beta=5\cdot 10^{-5}$. We note that our VAE is trained on $256\times 256$ images, and we also use it for our $512\times 512$ experiments without retraining (like [7]).
GIVT  For GIVT-Causal, we follow the original transformer decoder architecture [72] in decoder-only mode, but remove biases from attention layers, MLP blocks, and LayerNorms, and replace ReLU by GELU as is common practice. For GIVT-MaskGIT, we simply remove the attention mask during training and feed masked inputs instead of shifted ones. We use the BERT-Large configuration [13] by default (304M parameters), and also explore a larger backbone with 1.67B parameters, denoted with the suffix “-L” (see App. 0.B for details). For model variants with adapter (suffix “+A”), we use a stack of 8 bijective iRevNet blocks [28] (with hidden channel dimension $4d$, resulting in 112k additional parameters for $d=16$), applied to the $w\times h\times d$ representation before reshaping it into a sequence. We configure our GIVT models to predict a $16$-mixture GMM with factorized components (i.e. the mixture components are multivariate Gaussians with diagonal covariance), and explore predicting a single, multivariate Gaussian modeling the full covariance matrix of the tokens as an alternative. For the conditional generation experiments, we use a learned embedding which we prepend to the embedded token sequence. To train GIVT, we use Adam with a cosine schedule (500 epochs; with linear warmup of 50 epochs), set the learning rate and weight decay to $10^{-3}$ and $10^{-4}$, respectively, the optimizer $\beta_{2}$ parameter to $0.95$, the dropout probability to $0.2$ for GIVT-causal and $0.4$ for GIVT-MaskGIT, and the batch size to $8192$. We use the same data augmentation as during VAE training (see [7, 46]), and sample from the VAE encoder distribution for every batch (an additional source of randomness besides data augmentation).
We implement GIVT in JAX [4] and use distrax [12] to implement the cand compute the log-probabilities.
GIVT-MaskGIT inference  Following [7], we fix the number of inference steps to 16 and employ the cosine schedule (i.e. letting $r=i/16$ at step $i$, the fraction of masked tokens is given by $\cos(\pi/2r)$). We also sort tokens by likelihood at each step and sample using a “choice temperature” $t_{C}$.
Exploring the VAE latent space  To better understand the interplay between the feature dimension $d$, the KL regularization $\beta$, the reconstruction quality of the VAE, and the sampling quality of the GIVT, we train VAEs with latent dimension in $\{4,8,16,32\}$ and $\beta$ in $\{2.5\cdot 10^{-5},5\cdot 10^{-5},10^{-4},2\cdot 10^{-4}\}$ using the VAE-training setup described at the beginning of this section. For each of the resulting VAEs we train a GIVT-Causal with the smaller BERT-Base [13] dimensions and a range of values for the number of mixtures $k$.
Evaluation  For the VAEs we report “reconstruction FID”, the FID obtained when reconstructing the 50k ImageNet validation images. For our GIVT variants and baselines, we report the sampling FID [23] when sampling a balanced set of 50k images covering all ImageNet classes. In both cases, we rely on the well-established ADM TensorFlow Suite [14], which uses the entire ImageNet training set as a reference. Furthermore, we also report Precision and Recall [58]. Finally, we evaluate the representation learning capabilities by training a linear classifier on an average-pooled intermediate representation of unconditional GIVT-Causal as in prior work [8, 76] (see App. 0.F for details).
4.2 Panoptic segmentation and depth estimation
We build on the UViM framework [33], which uses a VQ-VAE to compress the label space of computer-vision dense-prediction tasks, and an encoder-decoder transformer taking the RGB image as an input and predicting the associated dense labels as discrete codes in the VQ-VAE latent space. Here, we replace the VQ-VAE with a $\beta$-VAE and use a GIVT encoder-decoder to model the continuous latent code. For the VAE, we use the same transformer-based autoencoder architecture (6-layer encoder and 12-layer decoder) and cross-entropy loss as [33]. We set $d=16$, $k=1$, and KL weight $\beta=2.5\cdot 10^{-4}$ for panoptic segmentation and $\beta=2\cdot 10^{-4}$ for depth estimation. To build an encoder-decoder GIVT model the same way as in [33], we employ the causal variant described for ImageNet generation and insert a cross-attention layer after each self-attention layer. Following [33], we use the ImageNet-21k-pretrained ViT-L/16 from [62] as the encoder, set the image resolution to 512px, and adopt the preprocessing and optimization hyper-parameters from [33]. We use the UViM variant without encoder and decoder context [33]. Finally, we consider variance scaling and beam search, selecting the parameters on a held-out subset of the training set as [33].
5 Results
Table 2: Results on class-conditional $512{\times}512$ ImageNet. We use the standard ADM evaluation suite for metrics, where FID is calculated w.r.t. the training set. GIVT-MaskGIT obtains competitive FID scores with only 16 inference steps and outperforms its VQ-counterpart. GIVT-Causal-L+A outperforms the best DiT variant, DiT-XL/2-G. Values obtained by us from public code. Inference uses activation caching.
[…]
ADM [14] 250 $\numprint{23.2}$0 $\numprint{0.73}$ $\numprint{0.60}$
ADM-G [14] CG $=1.0$ 250 $\numprint{7.72}$ $\numprint{0.87}$ $\numprint{0.42}$
DiT-XL/2 [51] 250 $\numprint{12.03}$ $\numprint{0.75}$ $\numprint{0.64}$
DiT-XL/2-G [51] CFG $=1.5$ 250 $\numprint{3.04}$ $\numprint{0.84}$ $\numprint{0.54}$